Grade 12Physical Chemistry

GIVE NACL REQUIRE 3 MOLES OF H20 AND ASKED CHARGE REQ FOR ELECTROLYSIS WITH 75% EFFICIENCY

Rajasai Bandaru
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
From the Faraday’s eectrolysis law,
moles x n-factor = charge/F
3*1 = q/F
q = 3F
with 75% efficiency chege becomes =>
q = 0.75*3F = > 9F/4 Coulamb where F =96500